Solve:
3x - y = 4
xy = 7

Answer by mukhopadhyay(490)

xy = 7
=>x=7/y
3x - y = 4
=>3(7/y)-y=4
=>21/y-y=4
=>21-y^2=4y
=>y^2+4y-21=0
=>y^2+7y-3y-21=0
=>y(y+7)-3(y+7)=0
=>(y+7)(y-3)=0
=>y=-7 or y=3
When y=-7, x=7/y=7/(-7)=-1
When y=3, x=7/3
Answer {(-1,-7),(7/3,3)}